# Cavity Control of Topological Qubits: Fusion Rule, Anyon Braiding and Majorana-Schrödinger Cat States

### v1

@@ Abstract
We investigate the impact of introducing a local cavity within the center of a topological chain, revealing profound effects on the system’s quantum states. Notably, the cavity induces a scissor-like effect that bisects the chain, liberating Majorana zero modes (MZMs) within the bulk. Our results demonstrate that this setup enables the observation of non-trivial fusion rules and braiding –key signatures of non-Abelian anyons– facilitated by the spatially selective ultra-strong coupling of the cavity photon field. These MZM characteristics can be directly probed through fermionic parity readouts and photon Berry phases, respectively. Furthermore, by leveraging the symmetry properties of fermion modes within a two-site cavity, we propose a novel method for generating MZM-polariton Schrödinger cat states. Our findings present a significant advancement in the control of topological quantum systems, offering new avenues for both fundamental research and potential quantum computing applications.
@@
Introduction.— Topological quantum computing is a holy grail of current research on quantum technologies [1]. Its key building blocks are non-abelian anyon quasi-particles in topological matter [2, 3, 4]. Clear experimental evidence on such anyons is still missing but recent advances in precision control of hybrid semiconductor-superconductor nanowires and quantum dot arrays have brought researchers closer to their realization [5, 6]. This opens new possibilities in designing novel experimental frameworks while making conceptual questions of great relevance.
@@
The essential physics of 1D Ising anyons is captured by the Kitaev chain (KC) model [7]. Each local Dirac fermion of the chain can be expressed in terms of a pair of Majorana fermions (MF). A key consequence is the existence of a topological phase in which a zero energy mode exists. It consists of two MF each one localized at one edge of the chain. Occupied and unoccupied options of such state have zero energy, forming a non-local qubit. Due to the splitting of the Majorana zero mode (MZM), quantum information is stored in both halves (data redundancy). If one half of the mode suffers interference, there is still enough information stored in the other half to proceed with the computation. Moreover, MZM become decoupled from the rest of the KC being an excellent candidate to present non-abelian statistics. Topological protection is produced through braiding and measurement of the non-abelian anyon character of MZM.
Recently, braiding and quantum teleportation of delocalized Majorana qubits have been established in KC manipulated by superconducting resonators [8, 9]. This remarkable experimental achievement prompts us to elaborate on extended scenarios to theoretically investigate for further non-local features in condensed matter systems aiming to put on firm grounds topological quantum computation routines. The most popular scheme proposed to achieve anyon braiding in topological 1D wire networks relies on a T-junction geometry [10, 11, 12]. The goal of most studies has been to explore how the pairing between different MF featuring an MZM evolves under a cyclic tuning of different coupling schemes. These simulations can be challenging both experimentally and theoretically. On the other side, the question of what might happen with an anyon fusion or braiding protocol relying on a purely geometric quantum state property has received far less attention.
The present work aims to pave the way to an alternative to multi-wire T-junction schemes: the analysis of the potential capabilities of a wire topological network controlled by the coupling to a bosonic environment in the regime of cavity quantum electrodynamics. This would open the use of photons and topological matter-connected states for new strategies to harness non-classical aspects of both matter and radiation ranging from fundamental physics aspects to potential applications in quantum information/computation issues. This becomes even more relevant in powerful platforms with novel quantum materials strongly coupled to optical and microwave cavities [13, 14, 15, 16, 17, 18] where photon quantum states, specifically coherent and squeezed states, show great promise in this regard due to the ease of availability and well-established control techniques.
@@
In stark contrast to previous proposals that aimed the demonstration of Majorana fusion rule using diverse connecting elements (quantum dots or electrostatic gates), to mediate the exchange between different KC edge MZMs [11, 12, 19, 20], our proposed KC-cavity set up is based on a single KC in the deep topological phase. In this way, the cavity acts as a scissor, splitting the original KC into two halves and consequently generating two new MZMs (for a fully electrostatic cutting way see [21]). The basic idea of previous works with cavity-controlled MZM systems [22, 23, 24] was primarily focused on the lowest-order photon-MZM coupling effects, i.e. where pre-defined MZM already existed before embedding the topological matter within the cavity, instead of a genuine full polariton image. Here, we propose a significantly distinct physics framework: the local cavity acts as a topological scissor [25], giving rise to a set of MZMs at the ends of the two parts of the original KC outside the cavity. Furthermore, it enables the simulation of their exchange by gathering up the precise Berry phase of the cavity field under an adiabatic cyclic variation of the matter-light coupling strength. As we will demonstrate below, for low and intermediate coupling strengths the polariton state is an entangled matter-light state. Nevertheless, in the ultra-strong coupling (USC) regime, the asymptotic decoupling of light and matter degrees of freedom [26, 27] enables both anyon fusing and braiding processes, mediated by photons, to occur in a KC-cavity system. Moreover, highly non-classical photon states in the cavity, Schrödinger cat states assisted by Majorana nonlocal qubits, become accessible. Such features turn out to be within reach in the USC regime due to the burgeoning interest on new cavity quantum electrodynamics (cQED) architectures and material supports [14, 15, 16]. The USC is particularly interesting for going beyond such first-order effects. The transition between different photon coherent states goes with corresponding modifications of fermion correlations. Moreover, of particular interest is the fact that in the asymptotic decoupling regime, cavity decoherence effects turn out to be innocuous for the MZM dynamics.
Model system.— We consider a single KC with vanishing chemical potential $\mu=0$, and identical hopping ($t$) and pairing ($\Delta$) terms, i.e. $t=\Delta$, the so-called sweet-spot for a quantum dot realization of the minimal KC [21]. For vanishing chemical potential the Kitaev Hamiltonian takes a diagonal form in terms of new non-local fermion operators $\hat{d}_{j}$ and ${\hat{d}_{j}^{\dagger}}$, i.e. $\hat{H}_{\rm KC}=2\Delta\sum_{j=1}^{N-1}\left[\hat{d}_{j}^{\dagger}\hat{d}_{j}-\frac{1}{2}\right]=i\Delta\sum_{j=1}^{N-1}\hat{\gamma}_{j,2}\hat{\gamma}_{j+1,1}$ (see Supplemental Material (SM) [*[SeetheSupplementalMaterialat][fordetailsofthecalculationsandderivations.]SM_Quiroga] and Ref. [29] for their definition in terms of both original Kitaev local operators ($\hat{c}_{j}^{\dagger},\hat{c}_{j}$) or Majorana fermions ($\hat{\gamma}_{j,1},\hat{\gamma}_{j,2}$)). The single mode microcavity Hamiltonian takes the usual form $\hat{H}_{\rm C}=\omega\hat{a}^{\dagger}\hat{a}$, where $\omega$ is the photon quantum energy ($\hbar=1$). Finally, the KC-cavity coupling term is given by $\hat{H}_{X}=\frac{1}{\sqrt{n_{\rm Cav}}}\sum_{j\in{\rm Cav}}\lambda_{j}\hat{c}_{j}^{\dagger}\hat{c}_{j}\left(e^{-i\phi_{j}}\hat{a}^{\dagger}+\hat{a}e^{i\phi_{j}}\right)$ with $\lambda_{j}$ the coupling strength between the KC $j$-site and the cavity field, ${\rm Cav}$ is the set with the site indices interacting with the cavity and $\phi_{j}$ controllable phases [30].
@@
Similarly, for a cavity enclosing two contiguous KC sites (from now on $2CK$ case, Fig. 1(b), sites $j=s$ and $j=s+1$, or equivalently only three $d$-fermion operators ($\hat{d}_{s-1},\hat{d}_{s}$ and $\hat{d}_{s+1}$) are brought into play so just $4$ matter states are involved in the full system quantum evolution, a two-qubit effective Rabi-like Hamiltonian arises:
$\begin{split}\hat{H}_{2}&=-\Delta\left(\hat{\sigma}_{z,1}+\hat{\sigma}_{z,2}+\hat{\sigma}_{z,1}\hat{\sigma}_{z,2}\right)+\omega\hat{a}^{{\dagger}}\hat{a}\\ &+\frac{1}{2\sqrt{2}}\sum_{r=1}^{2}\lambda_{r}\left(e^{-i\phi_{r}}\hat{a}^{\dagger}+e^{i\phi_{r}}\hat{a}\right)\left(1-\hat{\sigma}_{x,r}\right),\end{split}$ (2)
where $\hat{\sigma}_{\nu,r}$, with $\nu=x,y,z$, denote Pauli matrices for spin $r=1,2$, complex valued $\lambda_{1}e^{-i\phi_{1}}$ and $\lambda_{2}e^{-i\phi_{2}}$ represent the coupling strengths at sites $s$ and $s+1$, respectively (see SM [28]). Restriction to the even fermion parity sector has been kept. Notice that for both $1CK$ and $2CK$ cases, the resonance condition between matter and photonic sub-systems is $\omega=4\Delta$ and the $d$-fermion parity is a constant of motion, i.e. a single photon can only produce changes on the $d$-fermion occupation number by $0,\pm 2$.

### v2

@@ Abstract
We investigate the effects of coupling a local electromagnetic cavity to a segment of a topological Kitaev chain (KC), with particular emphasis on the interplay between photons and Majorana zero modes (MZMs). In addition to the well-known scissor effect-which effectively partitions the chain and isolates free MZMs in the bulk-we provide evidence of non-trivial fusion rules and braiding operations, hallmark signatures of non-Abelian anyons, enabled by spatially selective ultrastrong KC-cavity coupling. We propose that these distinctive MZM properties can be experimentally probed via fermionic parity measurements and photon-induced Berry phases. Furthermore, we demonstrate that, in the so-called sweet-spot regime, the coupled system can be mapped onto a Rabi-like model with a homodyne-rotated quadrature, offering a simplified yet powerful theoretical description. Exploiting the symmetry of fermionic modes within a two-site cavity configuration, we also show the feasibility of generating hybrid MZM-polariton Schrödinger cat states. Our findings offer a novel approach to manipulating topological quantum matter through local light-matter interactions and provide theoretical tools for future experimental realizations in platforms such as quantum materials or superconducting circuits.
: Phys. Scr. 100, 085122 (2025)
Keywords: Majorana Zero Modes, Kitaev Chain, non-Abelian anyons, Fusion Rule, Braiding
1 Introduction
@@
Topological quantum computing is the holy grail of current research on quantum technologies [1]. Its key building blocks are non-Abelian anyon quasi-particles in topological matter [2, 3, 4]. Clear experimental evidence on such anyons is still missing, but recent advances in precision control of hybrid semiconductor-superconductor nanowires and quantum dot arrays have brought researchers closer to their realization [5, 6]. This opens new possibilities in designing novel experimental frameworks while raising conceptual questions of great relevance.
The essential physics of 1D Ising anyons is captured by the Kitaev chain (KC) model [7]. In recent years, quantum-dot (QD) superconductor arrays have emerged as a promising platform for realizing Kitaev chains [8, 9]. Notably, a minimal two-site configuration has been successfully implemented in low-dimensional semiconductor systems [10], with tunneling spectroscopy providing direct evidence of MZM at a specific parameter regime [6, 11, 12, 13]. These advancements position hybrid semiconductor-superconductor heterostructures as highly versatile platforms, driving significant progress in developing topological solid-state qubits [14, 15]. A particularly promising configuration is the so-called “sweet spot” regime, where fine-tuned experimental conditions highlight Majorana-like physics, even though the system may lack full topological protection.
@@
In a KC, each local Dirac fermion of the chain can be expressed in terms of a pair of Majorana fermions (MF). A key consequence is the existence of a topological phase in which a zero-energy mode exists. It consists of two MF each one localized at one edge of the chain. Occupied and unoccupied options of such state have zero energy, forming a non-local qubit. Quantum information is encoded in the collective state of both Majoranas, and this nonlocal nature imparts redundancy and resilience to decoherence. Moreover, the zero-energy Majorana modes (MZMs) become decoupled from the rest of the KC, making them ideal candidates for non-Abelian statistics. Topological protection is produced through braiding and measurement of the non-Abelian anyon character of MZM.
Recent quantum simulation experiments have significantly advanced our understanding of braiding and quantum teleportation using delocalized Majorana qubits, particularly in Kitaev chains coupled to superconducting resonators [16, 17]. These developments motivate further theoretical frameworks aimed at capturing dynamic light-matter interactions and their role in non-local parity manipulation.
The most established approach to implementing anyon braiding in topological systems relies on multi-wire 2D networks using T-junction [18, 19, 20] or Y-junction geometries [21, 22, 23, 24]. These schemes typically involve dynamically tuning the coupling between spatially separated Majorana fermions, a process that can be technically demanding both experimentally and theoretically. In contrast, the potential for realizing fusion and braiding protocols through purely geometric or quantum state-based mechanisms—with no explicit spatial motion—has received relatively little attention.
In this work, we propose an alternative approach to conventional multi-wire architectures by investigating the properties of a single Kitaev chain locally coupled to a quantized bosonic mode within the framework of cavity quantum electrodynamics. Our aim is to explore Majorana fusion and non-Abelian features through photon-mediated interactions in one-dimensional systems, eliminating the need for multi-dimensional braiding geometries. We specifically focus on the “sweet spot” regime, which—despite lacking full topological protection—has proven experimentally suitable for generating Majorana-like excitations in minimal quantum-dot-superconductor configurations [8, 10, 6, 11, 12, 13]. These excitations, often called “poor man’s Majorana modes” (PMMMs), retain non-locality and parity coherence, providing a controlled setting for probing Majorana physics in tunable platforms.
We demonstrate that coupling a minimal Kitaev chain to a single-mode quantum cavity in the ultrastrong coupling (USC) regime enables photon-driven manipulation of non-local fermionic parity sectors. The cavity acts as a dynamical “quantum scissor” [25], effectively splitting the chain and releasing PMMMs into the bulk. This mechanism enables simulation of anyon fusion, emergence of parity observables, and formation of entangled Majorana–Schrödinger cat states.
Crucially, unlike previous works that treated MZMs as static inputs to a cavity system [26, 27, 28, 29], our approach treats the coupled light-matter system as a unified polaritonic entity. This viewpoint reveals Berry-phase-like signatures and parity control under adiabatic modulation of light–matter coupling, which can be engineered using superconducting circuits or hybrid semiconductor-superconductor devices [30, 31, 32].
Altogether, our findings establish a photonic framework to emulate and manipulate Majorana fusion physics in one-dimensional systems, broadening the landscape of quantum simulations of non-Abelian anyons and paving the way for light-assisted quantum operations in topologically inspired settings.
@@
As we will demonstrate below, for low and intermediate coupling strengths the polariton state is an entangled matter-light state. Nevertheless, in the ultra-strong coupling regime, the asymptotic decoupling of light and matter degrees of freedom [33, 34] enables both anyon fusing and braiding processes, mediated by photons, to occur in a KC-cavity system. Additionally, our proposal gives access to highly non-classical Schrödinger cat photon states, supported by the underlying nonlocal Majorana structure. Such features turn out to be within reach in the USC regime due to the burgeoning interest on new cavity quantum electrodynamics (cQED) architectures and material supports [35, 30, 31]. The USC is particularly interesting for going beyond such first-order effects. The transition between different photon coherent states goes with corresponding modifications of fermion correlations. Remarkably, in the asymptotic decoupling regime, cavity decoherence becomes harmless to MZM dynamics, allowing non-destructive readout of topological qubits. This provides a distinct advantage over existing QD-based approaches [36, 37]. Thus, achieving both the sweet-spot in Kitaev chains and the USC regime in cQED setups opens new routes for probing MZM statistics via light.
This paper is organized as follows: Section 2 briefly discusses the KC-(local) cavity model and its transformation to Rabi-like models in the so-called sweet-spot parameter regime as considered in this study as well as its theoretic background. In Section 3, we discuss the anyon fusion rule for a single site cavity. In Section 4, we present and discuss the braiding protocol for a two-site cavity. Next, in Section 5 Majorana-Schrödinger cat states are addressed. In Section 6, we summarize our findings, discuss potential future directions, and explore the possibilities for experimental validation. Further technical details have been collected in A and B.
2 Proposal for a photon assisted testbed of non-Abelian anyon properties
We consider a hybrid light-matter system consisting of a single-mode quantum cavity coupled to a Kitaev chain. The system is described by the Hamiltonian:
$\hat{H}=\hat{H}_{\rm C}+\hat{H}_{\rm KC}+\hat{H}_{\rm X}.$ (1)
Where, the single mode microcavity Hamiltonian takes the usual form $\hat{H}_{\rm C}=\omega\hat{a}^{\dagger}\hat{a}$, where $\omega$ is the photon quantum energy ($\hbar=1$), with $\hat{a}$($\hat{a}^{\dagger}$) the annihilation (creation) microwave photon operator. The Hamiltonian $\hat{\mathcal{H}}_{\rm KC}$ describing an isolated Kitaev chain with open boundaries takes the form [7]
$\hat{H}_{\rm KC}=-\frac{\mu}{2}\sum_{j=1}^{L}2(\hat{c}_{j}^{\dagger}\hat{c}_{j}-\hat{1})-\sum_{j=1}^{L-1}t(\hat{c}_{j}^{\dagger}\hat{c}_{j+1}+\hat{c}_{j+1}^{\dagger}\hat{c}_{j})-\Delta(\hat{c}_{j}\hat{c}_{j+1}+\hat{c}_{j+1}^{\dagger}\hat{c}_{j}^{\dagger}).$ (2)
Here, $\hat{c}_{j}$ and $\hat{c}_{j}^{\dagger}$ denote the fermionic annihilation and creation operators at lattice site $j=1,\ldots,L$, respectively. The parameter $\mu$ represents the chemical potential, $t$ is the nearest-neighbor hopping amplitude, and $\Delta$ characterizes the superconducting pairing strength between adjacent sites. This model describes a one-dimensional topological superconductor, featuring a quantum phase transition between topological and trivial (nontopological) phases at $\mu=2\Delta$, assuming $\Delta=t$.
One-dimensional topological superconductors are predicted to host robust Majorana zero modes [38]. However, in realistic implementations, microscopic disorder and finite-size effects can hinder the unambiguous detection and control of these modes [6, 39, 40]. An alternative approach that has gained increasing attention involves engineering effective Kitaev chains atom-by-atom [41] or through quantum dot arrays in hybrid semiconductor-superconductor systems [6, 42, 43].
Notably, even minimal systems—such as a two-site Kitaev chain-can support so-called “poor man’s Majorana modes” (PMMMs) when operated at the so-called “sweet spot”: vanishing chemical potential ($\mu=0$) and equal hopping and superconducting pairing amplitudes ($t=\Delta$[10]. While these states lack full topological protection, they retain key features of MZMs, including spatial separation and parity-based coherence [6].
Guided by both experimental feasibility and analytical tractability, we consider a Kitaev chain in this sweet spot regime as the basis for our photon-assisted architecture. This minimal configuration provides a highly controllable platform for exploring emergent non-Abelian features, especially when interfaced with quantized cavity fields [44, 42]. For vanishing chemical potential the Kitaev Hamiltonian takes a diagonal form in terms of new non-local fermion operators $\hat{d}_{j}$ and $\hat{d}_{j}^{\dagger}$}, i.e. $\hat{H}_{\rm KC}=2\Delta\sum_{j=1}^{N-1}\left[\hat{d}_{j}^{\dagger}\hat{d}_{j}-\frac{1}{2}\right]=i\Delta\sum_{j=1}^{N-1}\hat{\gamma}_{j,2}\hat{\gamma}_{j+1,1}$ (see Ref. [45] for their definition in terms of both original Kitaev local operators ($\hat{c}_{j}^{\dagger},\hat{c}_{j}$) or Majorana fermions ($\hat{\gamma}_{j,1},\hat{\gamma}_{j,2}$)).
Finally, the KC-cavity coupling term is given by
$\hat{H}_{X}=\frac{1}{\sqrt{n_{\rm Cav}}}\sum_{j\in{\rm Cav}}\lambda_{j}\hat{c}_{j}^{\dagger}\hat{c}_{j}\left(e^{-i\phi_{j}}\hat{a}^{\dagger}+\hat{a}e^{i\phi_{j}}\right),$ (3)
where $\lambda_{j}e^{-i\phi_{j}}$ the complex-valued coupling strength between the KC $j$-site and the cavity field, Cav is the set with the site indices interacting with the cavity [46]. For convenience, we refer to a single Kitaev site coupled to the cavity as $1CK$, and to two Kitaev sites coupled to the cavity as $2CK$ throughout this work.
2.1 Bulk single site case
We consider the case when the chain-cavity coupling contribution at site $j=s$ of the KC in Eq. (3) is the fermion number operator $\hat{c}_{s}^{\dagger}\hat{c}_{s}=\frac{1}{2}\left(1+i\hat{\gamma}_{s,1}\hat{\gamma}_{s,2}\right)=\frac{1}{2}\left(1+\hat{d}_{s-1}\hat{d}_{s}+\hat{d}_{s-1}\hat{d}_{s}^{\dagger}-\hat{d}_{s-1}^{\dagger}\hat{d}_{s}-\hat{d}_{s-1}^{\dagger}\hat{d}_{s}^{\dagger}\right)$. In panel (a) of Fig. 1, we schematically represent the bulk single-site case. Remarkably, for this coupling term the parity of the number of $d$-fermions is a constant of motion, i.e. a single photon can only produce changes on the $d$-fermion number by $0,\pm 2$. Since only two $d$-fermion operators ($\hat{d}_{s-1}$ and $\hat{d}_{s}$) are included in the matter-radiation coupling term, just $4$ matter (KC) states are involved in the full system quantum evolution. Therefore, the Kitaev even state sector is spanned by the basis set $|-\rangle_{z}=|{\rm vac}\rangle=|\circ\circ\rangle$, $|+\rangle_{z}=\hat{d}_{s-1}^{\dagger}\hat{d}_{s}^{\dagger}|\circ\circ\rangle=|\bullet\bullet\rangle$. Where $|{\rm vac}\rangle=|\circ\circ\rangle$ represents the vacuum state, i.e. no $d$-fermion excitations or $|\circ\circ\rangle=0$ for j=s1,sj=s-1,s. Consequently, the $1CK$ Hamiltonian reduces to that of an effective Rabi-like model with homodyne rotated quadrature:
We proceed to obtain an approximate energy spectrum in the USC regime where the first term in Eq. (4) is treated as a perturbation. At first order in $\Delta$, two-dimensional subspaces, spanned by basis sets $\{|-\rangle_{x}\otimes\hat{D}(\alpha)|n\rangle,|+\rangle_{x}\otimes|n^{\prime}\rangle\}$, $n,n^{\prime}=0,1,2,\ldots$, with $\alpha=-\frac{\lambda}{\omega}e^{-i\phi}$ and $\hat{D}(\alpha)$ the displacement operator, are separately diagonalized yielding to:
$E_{n,\pm}=\frac{1}{2}\left[2n\omega-\frac{\lambda^{2}}{\omega}\pm\sqrt{\frac{\lambda^{4}}{\omega^{2}}+16\Delta^{2}e^{-|\alpha|^{2}}L_{n}^{2}\left[|\alpha|^{2}\right]}\;\right],$ (5)
where $L_{n}(x)$ is a Laguerre polynomial. In the weak-coupling regime ($\lambda<\omega\sim\Delta$) matter states are suited best for exploring the dynamics controlled by $\hat{H}_{1}$. However, in the USC regime ($\lambda>\omega\sim\Delta$) a better-adapted basis set is provided by occupation number states, $|n_{L,R},n_{c}\rangle$, of a non-local fermion $f_{L,R}=\frac{1}{2}\left(\hat{\gamma}_{L}+i\hat{\gamma}_{R}\right)$ and a local fermion inside the cavity, $c_{s}=\frac{1}{2}\left(\hat{\gamma}_{s,1}+i\hat{\gamma}_{s,2}\right)$. The basis transformation is:
$|\circ\circ\rangle$ $=\frac{1}{\sqrt{2}}\left(|\circ_{L,R},\circ_{c}\rangle-i|\bullet_{L,R},\bullet_{c}\rangle\right),$
$|\bullet\bullet\rangle$ $=\frac{1}{\sqrt{2}}\left(|\circ_{L,R},\circ_{c}\rangle+i|\bullet_{L,R},\bullet_{c}\rangle\right).$ (6)
In this regime the ground state is simply expressed as $|GS\rangle\simeq-i|1_{L,R},1_{c}\rangle\otimes|\alpha\rangle$ with $\alpha=-\frac{\lambda}{\omega}e^{-i\phi}$.
2.2 Two bulk sites case
The cavity couples with $2$ nearest-neighbor bulk physical sites ($s$ and $s+1$ with s=2,,N2s=2,...,N-2), or equivalently $3$ bulk bond fermion modes are involved) characterized by the interaction Hamiltonian:
@@ 2.2 Two bulk sites case
where the matter operator $\hat{h}_{r}$ is given by $1+\hat{d}_{s+r-1}\hat{d}_{s+r}+\hat{d}_{s+r-1}\hat{d}_{s+r}^{\dagger}-\hat{d}_{s+r-1}^{\dagger}\hat{d}_{s+r}-\hat{d}_{s+r-1}^{\dagger}\hat{d}_{s+r}^{\dagger}$. Where complex valued $\lambda_{1}e^{-i\phi_{1}}$ ($\lambda_{2}e^{-i\phi_{2}}$) denote the coupling strength at site $s$ ($s+1$). The total Hamiltonian conserves the parity of the number of fermions, i.e. it can only produce changes on the fermion number by $0,\pm 2$. Given the fact that the non-interacting matter-field ground-state has zero $d$-fermions, jointly with zero photons, it is natural to confine this study to the even sector of the fermion Hilbert space. Thus, just $4$ matter (KC) states are involved in the full system quantum evolution. A mapping of the matter states to a two-spin system is thus immediate. In the following, we restrict ourselves to the even KC-sector and $\phi_{1}=0$. For $\phi_{2}$, two cases will be addressed, $\phi_{2}=0$ ($\lambda_{2}>0$) or $\phi_{2}=\pi$ ($\lambda_{2}<0$). In a two-spin basis formed by eigenvectors of $z$-Pauli matrix components $\{|+,+\rangle_{z},|+,-\rangle_{z},|-,+\rangle_{z},|-,-\rangle_{z}\}$, the KC states are now associated to:
$|-,-\rangle_{z}$ $=|\circ_{s-1},\circ_{s},\circ_{s+1}\rangle,\quad|-,+\rangle_{z}=|\circ_{s-1},\bullet_{s},\bullet_{s+1}\rangle,$
$|+,-\rangle_{z}$ $=|\bullet_{s-1},\bullet_{s},\circ_{s+1}\rangle,\quad|+,+\rangle_{z}=|\bullet_{s-1},\circ_{s},\bullet_{s+1}\rangle.$ (8)
The free KC term in the Hamiltonian acquires an Ising-type spin-spin coupling term, yielding finally a two-qubit effective Rabi-like Hamiltonian:
$\hat{H}_{2}=$ $-\Delta\left(\hat{\sigma}_{z,1}+\hat{\sigma}_{z,2}+\hat{\sigma}_{z,1}\hat{\sigma}_{z,2}\right)+\omega\hat{a}^{{\dagger}}\hat{a}$ (9)
$+\frac{1}{2\sqrt{2}}\sum_{r=1}^{2}\lambda_{r}\left(e^{-i\phi_{r}}\hat{a}^{\dagger}+e^{i\phi_{r}}\hat{a}\right)\left(1-\hat{\sigma}_{x,r}\right),$
where $\hat{\sigma}_{\nu,r}$, with $\nu=x,y,z$, denote Pauli matrices for spin $r=1,2$, complex valued $\lambda_{1}e^{-i\phi_{1}}$ and $\lambda_{2}e^{-i\phi_{2}}$ represent the coupling strengths at sites $s$ and $s+1$, respectively (see A). Restriction to the even fermion parity sector has been kept. Notice that for both $1CK$ and $2CK$ cases, the resonance condition between matter and photonic sub-systems is $\omega=4\Delta$ and the $d$-fermion parity is a constant of motion, i.e. a single photon can only produce changes on the $d$-fermion occupation number by $0,\pm 2$.
On the other hand, the convenient and novel Hamiltonian structure, described as a single-qubit or two-qubit Rabi-like model for the $1CK$ and $2CK$ cases, respectively (as given by Eqs. (4) and (9)), facilitates both analytical and numerical advancements. In quantum optics, the USC regime of light-matter interactions has been the subject of intense investigation from both theoretical and experimental perspectives [47]. In particular, the role of the diamagnetic term in the USC regime has attracted significant attention [48, 49, 50, 51] and criticism [52, 53]. For this study, we adopt this framework as a complete and valid description for the USC regime.